OpenStudy (anonymous):

Help Please!!! Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. (5 points) Part B: How many liters of orange juice are present in the can? Show your work. (5 points)

OpenStudy (mathstudent55):

The total content of the can is x. That means that part of x is orange juice and part of x is water.

OpenStudy (anonymous):

Okay.., how would i make this a function. I am no could with these! :(

OpenStudy (mathstudent55):

Since the can has only orange juice and water and the orange juice part is 80%, that means the water part must be 20%. Ok so far?

OpenStudy (anonymous):

20% of x = 1 liter?

OpenStudy (mathstudent55):

That means the equation for the water part is: 20%x = 1

OpenStudy (mathstudent55):

Wow, you beat me to it. Good job!

OpenStudy (anonymous):

so there are 5 liters total? 4 liters = orange juice, 1= water

OpenStudy (mathstudent55):

For part B, first solve for x, the total amount of liquid in the can.

OpenStudy (anonymous):

i mean equations LOL

OpenStudy (mathstudent55):

0.2x = 1 x = 5 There are 5 liters of liquid Orange juice = 80%x = 0.8 * 5 = 4 liters.

OpenStudy (mathstudent55):

You keep getting the answers before me. I think I should be asking for your help instead of you asking me for help!

OpenStudy (anonymous):

Haha. SO what would be the equations for part 1? I can never seem how to put the information in a function

OpenStudy (anonymous):

I MEAN EQUATION

OpenStudy (anonymous):

not function..

OpenStudy (anonymous):

I need help with the equation for part a? how do i form. @mathstudent55

OpenStudy (anonymous):

0.2x = 1 x = 5 There are 5 liters of liquid Orange juice = 80%x = 0.8 * 5 = 4 liters. was that it? LOL

OpenStudy (mathstudent55):

We did the equation above: 20%x = 1 or simply 0.2x = 1

OpenStudy (anonymous):

THANK YOU :D